# Beyond Grids: Multi-objective Bayesian Optimization With Adaptive Discretization

### v3

@@ A.7 Derivation of the dimension-type sample complexity bound
$\overline{\omega}^{2}_{t}$ $=\omega^{2}_{t}(x_{h_{t},i_{t}})=\Big{(}\max_{y,y^{\prime}\in\bm{R}_{t}(x_{h_{t},i_{t}})}||y-y^{\prime}||_{2}\Big{)}^{2}$
$\leq\Big{(}\max_{y,y^{\prime}\in\bm{Q}_{t}(x_{h_{t},i_{t}})}||y-y^{\prime}||_{2}\Big{)}^{2}=\Big{(}||\bm{U}_{t}(x_{h_{t},i_{t}})-\bm{L}_{t}(x_{h_{t},i_{t}})||_{2}\Big{)}^{2}=\sum^{m}_{j=1}\left(\overline{B}^{j}_{t}(x_{h_{t},i_{t}})-\underline{B}^{j}_{t}(x_{h_{t},i_{t}})+2V_{h_{t}}\right)^{2}$
@@
Let ${\cal E}$ denote the set of rounds in which Adaptive $\bm{\epsilon}$-PAL performs design evaluations. Note that $|{\cal E}|=\tau_{s}$. Let $t_{s}$ denote the round in which the algorithm terminates. Next, we use Lemma 1 together with Lemma 14 to upper bound $\overline{\omega}_{t_{s}}$ as a function of the number evaluations until termination.
@@
$K_{1}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{-(\bar{D}+\alpha)}{\bar{D}+2\alpha}}+K_{2}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{\alpha}{\bar{D}+2\alpha}}\leq\epsilon~{},$
@@
Proof. According to Lemma 3, we have $\overline{w}_{t_{s}}\leq\epsilon$. In addition, Lemma 15 says that
@@
$T=\min\left\{\tau\in\mathbb{N}:K_{1}\tau^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log\tau\right)^{\frac{-(\bar{D}+\alpha)}{\bar{D}+2\alpha}}+K_{2}\tau^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log\tau\right)^{\frac{\alpha}{\bar{D}+2\alpha}}\leq\epsilon\right\}~{}.$
Since the event $\{\tau_{s}=T\}$ implies that $\{\overline{w}_{t_{s}}\leq\epsilon\}$, we have $\Pr(\tau_{s}>T)=0$.∎
A.8 The final step of the proof of Theorem 1
@@ A.9 Proof of Proposition 2
A.10 Determination of $\alpha$ and $C_{K}$ for Squared Exponential type kernels
@@
Appendix B The proof of Proposition 1

### v4

@@ B.7 Derivation of the dimension-type sample complexity bound
[…]
@@
[…]
@@
$K_{1}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{-(\bar{D}+\alpha)}{\bar{D}+2\alpha}}+K_{2}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{\alpha}{\bar{D}+2\alpha}}<\epsilon~,$
@@
Proof. According to Lemma 3, we have $\overline{w}_{t_{s}}<\epsilon$. In addition, Lemma 15 says that
@@
$T=\min\left\{\tau\in\mathbb{N}:K_{1}\tau^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log\tau\right)^{\frac{-(\bar{D}+\alpha)}{\bar{D}+2\alpha}}+K_{2}\tau^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log\tau\right)^{\frac{\alpha}{\bar{D}+2\alpha}}<\epsilon\right\}~.$
Since the event $\{\tau_{s}=T\}$ implies that $\{\overline{w}_{t_{s}}<\epsilon\}$, we have $\Pr(\tau_{s}>T)=0$.∎
B.8 The final step of the proof of Theorem 1
@@ B.9 Proof of Proposition 2
B.10 Determination of $\alpha$ and $C_{K}$ for Squared Exponential type kernels
@@
Appendix C The proof of Proposition 1